# First Observation of Antiproton Annihilation At Rest on Argon in the LArIAT Experiment

### v1

@@ Abstract
We report the first observation and measurement of antiproton annihilation at rest on argon using the LArIAT experiment. Antiprotons from a charged particle test beam that come to rest inside LArIAT’s liquid argon time projection chamber (LArTPC) are identified through beamline instrumentation and LArTPC track reconstruction algorithms. The multiplicity of charged particle tracks originating from the annihilation vertex is manually assessed through hand-scanning, resulting in a distribution with a mean of 3.2 ± 0.4 tracks and a standard deviation of 1.3 tracks. This is consistent with an automated track reconstruction, which produces a mean of 2.8 ± 0.4 tracks and a standard deviation of 1.2 tracks. Good agreement is found between data and Monte Carlo simulations for both methods. Additionally, we report the shower multiplicity and particle identification of outgoing tracks, both of which align closely with theoretical predictions. These findings will contribute to modeling of intranuclear annihilation interactions on argon, including scenarios such as neutron-antineutron oscillations.
@@ I Introduction
Due to their distinct “star-shaped” final-state topology, antiproton () annihilations at rest have produced some of the most striking images in particle physics. First observed in 1955 at the Bevatron using photographic emulsion [1, 2, 3], annihilation on nucleons () have now been observed by a number of experiments, and on various target nuclei [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16].
The annihilation takes place preferentially at the surface of a nucleus [17] and results in a set of particles, primarily pions (), emanating isotropically from the annihilation vertex. The exact number and mixture of produced particle types varies depending on the target nucleus. Since a fraction of the pions from the annihilation have to travel through the nucleus before becoming visible in the detector, they can undergo a number of final-state interactions (FSI) processes: absorption ( or ), neutral pion production (), quasi-elastic scattering (), or single-charge exchange ( or ). These processes modify the kinematics of the pions or even prevent pions from exiting the nucleus entirely.
Because of these FSI, final-states of annihilation can be used as a proxy for that of theorized neutron-antineutron (-) oscillations [18] where the newly-converted immediately annihilates with a neighboring nucleon. Observing this baryon number violating process would satisfy one of the Sakharov conditions [19] and would have profound implications for theories on baryogenesis, the origin of matter-antimatter asymmetry, and large extra dimensions [20, 21, 22, 23]. Assuming isospin symmetry, and the similarity of protons and neutrons in terms of quark content, it can be assumed that annihilation interactions result in outgoing products that are largely similar in particle makeup and kinematic distribution to those that would emerge from a annihilation [24]. Due to the difficulty in performing antineutron scattering measurements and their resulting scarcity [25], annihilation are critical inputs for - oscillation models. More complete models, especially of FSI, are necessary to test automated reconstruction on these isotropic events to improve the sensitivity of future searches for - in large liquid argon time projection chambers (LArTPCs) like DUNE.
This paper presents the first-ever measurement of antiproton annihilation at rest on argon, performed with the liquid argon in a testbeam (LArIAT) experiment [26]. Antiproton are identified amongst the particles in the beamline, and a subset of these are identified as annihilating at rest inside a LArTPC. Their outgoing particles are compared to a dedicated Monte Carlo (MC) simulation of true antiprotons annihilating at rest in the LArTPC. The results presented here represent a first validation of the modeling of nucleon annihilation on argon and may be useful in constraining systematic uncertainties of FSI for pions produced by nucleon annihilation on argon.
@@ II The LArIAT Experiment
The LArIAT experiment ran at Fermilab’s Test Beam Facility [27] situated in the Meson Center beam line. LArIAT’s primary purpose has measuring charged particle interaction cross sections and developing and calibration techniques. The LArIAT TPC operated in conjunction with a series of beamline monitoring subsystems to enable the triggering of detector readouts for individual particles that traverse the entire beamline and interact within the chamber.
@@
Before reaching LArIAT, the secondary beam entered the experimental hall enclosure where it struck a copper target inside a steel collimator to produce a tertiary beam of charged particles. The tertiary beam traveled through a series of beamline instrumentation elements as shown in Fig. 1: two (upstream, downstream) beam collimators (USCOLL, DSCOLL), two (upstream, downstream) time-of-flight (TOF) scintillating paddles (USTOF, DSTOF), four multi-wire proportional chambers (MWPCs), two bending dipole magnets (B1, B2), and a HALO scintillating paddle used to identify and veto particles entering the LArTPC that did not follow the path of the tertiary beamline. The trajectory of a charged particle traveling in the beamline was reconstructed by interpolating the straight segments obtained by at least 3 hits in the upstream MWPC (WC1 and WC2 in Fig. 1) in tandem with the downstream ones (WC3 and WC4 in Fig. 1). This reconstructs the particle’s trajectory before and after the bending magnets which bends it in the -plane. The polarity of the magnets can be configured to select positively or negatively charged particles from the beam. The bending angle of the particle, together with the known magnetic field, are used to determine the particle’s momentum. The reconstructed beamline momentum for all triggered beamline particles in the dataset used for this analysis is shown in Fig. 2 using the A magnet current. In addition, scintillator paddles are used to determine the particle’s TOF, which is used in tandem with the momentum to determine the particle mass as explained in Sec. III. Detailed discussion on the TOF and MWPC systems can be found in [26, 28].
Tertiary beamline particles were then directed into the LArTPC, a 170-liter active LAr volume approximately 90 cm in length (, along the beam direction), 47 cm in width (), and 40 cm in height (). The LArTPC and cryostat are shown in Fig. 3. The reference coordinate system used in LArIAT originates from the midpoint of the anode plane of the LArTPC. The z-direction aligns with the beam line, y-direction points upward, and x-direction extends from the anode to the cathode. The LArTPC readout comprised two readout-instrumented wire planes composed of 240 wires each (induction and collection), as well as a non-instrumented 225-wire shield plane. For the data-taking campaign used in this study, wires were spaced 4 mm apart with a 4 mm separation gap between each of the planes. The induction and collection wire planes were oriented at 60° from the beamline direction. Ionization charge drifted along the direction from the cathode to the wire planes (anode) along a nominal field of 484 V/cm, which resulted in a maximum electron drift period of approximately 320 s. The LArIAT data acquisition system (DAQ) recorded the ionization produced by beamline particles traversing the LArTPC via induced or collected charge on the wire planes and these signals were digitized with a sampling period of 128 ns.
The analysis presented here uses beam-triggered data taken during 2016, referred to as Run II, when the tertiary beamline’s dipole magnets were operating with negative polarity in order to steer only negatively charged particles toward the LArTPC.
@@ III Beamline Reconstruction and Antiproton Selection
All beamline triggered events are reconstructed to obtain their individual TOF and momentum as they traverse the beamline monitoring systems. It is possible to distinguish different particle types by utilizing the reconstructed TOF and momentum. This is shown in Fig. 4 with the overlapping expectations for particles present in the beam relevant for this analysis: the black solid curve is the expected deuteron region, in red the proton region, and in orange the kaon region. The dense region around 23 ns is where all lower minimally ionizing particles are expected to be such as pions, muons, and electrons in green, blue, and purple, respectively. Faint horizontal bands present at TOF intervals of 20 ns are caused by multiple high-energy muons produced in the secondary beam far upstream of our tertiary beamline, with a time separation defined by the primary proton beam structure, which subsequently stream forward and separately strike the upstream and downstream TOF scintillator paddles creating a false TOF measurement for particular beam-triggered events.
@@
where is the speed of light and is the length traveled by the particle between the two TOF paddles, 6.65 m. The results for Run II negative polarity data for the calculation of the beamline masses are shown in Fig. 5. Particle types are classified according to three distinct regions as shown in the figure:
@@
The beamline selection cuts chosen for this analysis are tuned using the Run II positive polarity data where there exists a large sample of protons (100,000). A selection using the reconstructed mass between 800 MeV/c2 and 1100 MeV/c2, inclusively, is chosen to include all of the proton-like mass peak while minimizing contamination from kaons at lower mass and deuterons at higher mass. The selection is then applied to the data in Run II with -100 A magnet current used in this analysis to select the antiprotons. The impact of this selection cut is evident in the TOF and reconstructed beamline momentum distributions, as shown in Fig. 4, with the dashed curves the selected region for the antiprotons. Out of the 189,000 events processed in the Run II negative polarity dataset, 520 are selected as having a mass consistent with an antiproton. The region selected is shown by the red shaded region in Fig. 5.
Kaons could mimic incoming beamline candidates in the case of short beam particle track length due to their similar energy deposition profile. To assess potential kaon contamination in the sample, a combined fit was applied to the negative polarity data from Fig. 5, with a Gaussian function centered at the kaon mass of 494 MeV/c2 on top of a Landau function to model the non-kaon beam related backgrounds in the calculated mass distribution. These backgrounds fitted by the Landau are actually accidental TOFs from 20 ns-space halo muons striking the TOF paddles. Since the events in these horizontal bands are due to random firings of the trigger from muons produced upstream of the tertiary beamline, they feature a high number of tracks in the TPC with no clear stopping beam-matched track. These are unlikely to contaminate the sample and are removed during the LArTPC selection stage. Based on an extrapolation of the fit into the proton-mass selection window, expected contamination of kaons is highly negligible, with events expected for the full Run II sample.
@@ IV LArTPC Reconstruction and At-Rest Annihilation Selection
Following the antiproton beamline reconstruction and selection of Sec. III, the second reconstruction and selection phase happens in the LArTPC. In this phase, LArTPC data analysis employs a hybrid manual/automated reconstruction process, culminating in the determination of final candidates on an event-by-event basis using hand-scanning and hand-clustering methods as well as utilizing a ranged-based particle identification (PID) algorithm on both the incoming antiprotons and the outgoing particles from the annihilation vertex.
@@ IV.1 LArTPC Reconstruction & Simulation
LArTPC reconstruction is performed using the LArSoft (version 06.61.00) event processing framework [29, 26, 28] with outputs reconstructed particle objects and their associated kinematics. Particles passing through the LArTPC are reconstructed from ionization signals detected on the wires. These signals are waveforms of ADCs vs. time ticks; so-called “hits” within the waveforms are determined by Gaussian fits to over-threshold pulses found in filtered and deconvolved wire waveforms. Hits are typically clustered together by an automated reconstruction algorithm which reconstructs 3D tracks by time-matching clusters of hits across the two 2D wire-planes [30]. LArIAT’s automated reconstruction algorithms are optimized for forward-going particles and perform poorly on isotropic candidates by not being able to reconstruct clear and isolated particle tracks from various events. To improve performance, two manual components to the reconstruction are introduced in this analysis.
Hits associated with distinct particles are manually selected by grouping them together into 2D clusters on each wire plane in a customized event display graphical user interface. These clusters are then used to reconstruct 3D tracks using the standard projection-matching track fitting algorithm in LArSoft [31, 32]. Hits located directly at the annihilation vertex are excluded from clusters associated with outgoing particles to avoid complications in the tracking algorithm. This ensures that hits from the stopping antiproton’s Bragg peak or from overlapping outgoing ionization trails are not mistakenly included in the wrong clusters. For candidate incoming antiproton tracks, hand-clustering is treated more inclusively by extending the clustering process as close to the annihilation vertex as possible. An example of the results of individual and manual clustering is shown in Fig. 6 where the different clustered objects are individually grouped in different colors and noted from C1 through C5.
@@
In order to gauge the performance of the reconstruction methods here, a sample of 1,200 antiprotons entering the LArTPC is simulated. They are generated at the final beamline MPWC approximately 100 cm upstream of the TPC, with a momentum distribution roughly matching data in Fig 2. The MC simulation is performed this way to increase the number of annihilations at rest in the active area of the LArTPC mimicking reality as much as possible. The generated particles decay, interact, and deposit their energy following Geant4 [33] (version 10.3.1) using the QGSP_BERT physics library. To simulate antiproton annihilation on individual nucleons in argon, the Fritiof (FTF) [34, 35] model is used in this paper.
@@ IV.2 LArTPC Particle Identification and Event Selection
A clear, non-overlapping, incoming beamline particle track must come to a stop with a visible Bragg peak.
@@
Outgoing particles must start at the end of the incoming beamline particle and travel away from the annihilation vertex.
@@
The first criterion removes events where the annihilation might have occurred before entering the active region of the LArTPC resulting, in some cases, an empty event. In data, this criterion additionally removes events containing significant pile-up muon activity which may produce ionization that overlaps and interferes with the actual triggered antiproton track. These pile-ups are caused by additional tracks from muons produced in the upstream beamline elements traveling alongside the true beamline candidates. Since annihilation at rest produces outgoing particles isotropically which are required to be observed from the second criterion, it is possible that tracks in the LArTPC overlap with the incoming stopping antiproton. The clear Bragg peak of the selected events ensures that the particle is coming to rest with its track information passed to the d/d-based PID algorithm to identify the particle type.
@@
The particle type associated with a reconstructed track inside the LArTPC is identified by examining the calorimetry information for the track’s reconstructed hits. This is done by looking at the amount of energy deposited per unit length (d/d) along the 3D track object and comparing it to the expected values of d/d for different particle types. In this analysis, particles can be grouped as: proton-like, kaon-like, pion/muon-like, and minimum-ionizing particles without a Bragg peak (non-stopping-like). The PID method consists of calculating a between a track’s modified residual range profile and expected d/d profiles for different particle types, with the number of degrees of freedom (NDF) being the number of reconstructed hits used to calculate the total . This algorithm’s performance heavily depends on a well-defined and clear Bragg peak. Thus reconstructed hits directly at the annihilation vertex are omitted. This reduces tracking algorithm ambiguity due to charge from the stopping antiproton’s Bragg peak or with hits from overlapping outgoing ionization trails. The PID method utilized also accounts for the potentially shorter track length due to under clustering. The values of the residual range are corrected up to the largest expected uncertainty in the reconstructed annihilation vertex position (two wire spacings).
@@
After applying the selection criteria to the 1,200 simulated antiproton events, 80 stopping candidates are selected. By looking at the MC truth information: 69 are true annihilations at rest, while the other 11 retain some kinetic energy at the moment of annihilation. Among these, 9 events have a true final kinetic energy less than 60 MeV and are considered quasi-rest candidates. The two remaining events, each with nearly 200 MeV of remaining true kinetic energy, cannot be classified as quasi-rest candidates. Based on the obtained number of misidentified candidates, the mis-ID rate of the hand-scan method is estimated to be 6%. Of the initial 1,200 events, 130 are considered as signals by having true final kinetic energy of less than 60 MeV resulting in a selection efficiency of 60% by considering annihilation at rest and quasi-rest together.
@@
All of the 80 MC signal candidates are also processed through the PID algorithm. The results indicate that 70 of the 80 events are identified as proton-like, 6 as kaon-like, and 4 as stopping muon/pion-like. To evaluate the performance of the PID algorithm, the d/d distributions for the three categories of antiproton candidates are compared with the expected values, as shown in top of Fig. 7. The proton-like candidates closely align with the proton expectation curve. The kaon-like candidates are positioned between the proton and kaon expectations, leading the PID to favor the kaon hypothesis due to a slight bias towards this behavior. The stopping muon/pion-like events correspond well with the pion/muon expectation curve, thus can not be selected as antiprotons annihilating at rest. The two antiproton candidates that have over 200 MeV of kinetic energy are identified as 1 kaon-like and 1 muon/pion-like. Based on these observations, all 10 non-proton-like candidates are excluded from the sample resulting in a PID algorithm efficiency of 90%.
@@
Following the established selection criteria, the number of data events is reduced from 520 initial candidates in the selected mass range, down to 11 stopping antiprotons annihilation at rest candidate. All 11 candidates are identified as proton-like using the PID calculation. For completeness, the d/d as a function of the residual range distribution of the antiproton candidates is shown in bottom of Fig. 7 and all align well with the proton-like expected curve. All 11 candidates are expected to be true antiproton annihilations at rest given the relatively small mis-identification rate evaluated from MC. Furthermore, the distribution of the reconstructed beamline mass of these candidates is centered around the antiproton mass peak shown in Fig. 8 with a fitted Gaussian centered at 933 MeV/c2 enforcing the case that these are true antiprotons.
A summary of the effect of each selection cut is shown in Table 1 with the first line corresponding to the beamline selection for data, and all of the MC generated events. The second and third line corresponds to the hand-scanned selection while the fourth and final line to the PID using the LArTPC reconstructed hits. Pile-ups are not simulated in this analysis thus a difference in the relative number of events removed from data and MC is expected for the second selection cut.
@@
A collection of representative selected antiproton annihilation at rest event displays from data and MC simulation is provided in Appendix A.

### v2

@@ Abstract
We report the first observation and measurement of antiproton annihilation at rest on argon track and shower multiplicities and particle identification conducted with the LArIAT experiment. Stopping antiprotons from the Fermilab Test Beam Facility’s charged particle testbeam are identified using beamline instrumentation and LArIAT’s liquid argon time projection chamber (LArTPC). The charged particle multiplicity from the annihilation vertex is manually evaluated via hand-scanning, yielding a mean of 3.2 0.4 tracks and a standard deviation of 1.3 tracks, consistent with a semi-automated reconstruction resulting in 2.8 0.4 tracks and a standard deviation of 1.2 tracks. Both methods are consistent with Monte Carlo simulations within statistical uncertainty. The shower multiplicities and particle identification for outgoing tracks are also consistent with Geant4 models predictions. These results, obtained from a low-statistics sample, provide a foundation for higher-statistics studies in larger LArTPCs, which could refine modeling of intranuclear annihilation on argon and inform scenarios such as neutron-antineutron oscillations.
@@ I Introduction
Due to their distinct “star-shaped” final-state topology, antiproton () annihilations at rest have produced some of the most striking images in particle physics. First observed in 1955 at the Bevatron using photographic emulsion [1, 2, 3], annihilation on nucleons () has been observed by a number of experiments, and on various target nuclei [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16].
The annihilation takes place preferentially at the surface of the nucleus [17] and results in a set of particles, primarily pions (), emanating isotropically from the annihilation vertex. The exact number and mixture of produced particle types varies depending on the target nucleus and its density. Since a fraction [17, 18, 19] of the pions from the annihilation have to travel through the nucleus before becoming visible in the detector, they can undergo a number of final-state interactions (FSI) processes. These processes modify the kinematics of the pions or can prevent pions from exiting the nucleus entirely.
Assuming isospin symmetry, and the similarity of protons and neutrons in terms of quark content, it can be assumed that annihilation interactions result in outgoing products that are largely similar in particle makeup and kinematic distribution to those that would emerge from a annihilation [20]. Because of this, annihilation can be used as a proxy for that of theorized neutron-antineutron (-) oscillations [21] where the newly-converted immediately annihilates with a neighboring nucleon. Observing this baryon number violating process would satisfy one of the Sakharov conditions [22] and would have profound implications for theories on baryogenesis, the origin of the matter-antimatter asymmetry, and large extra dimensions [23, 24, 25, 26]. Due to the difficulty in performing antineutron scattering measurements and their resulting scarcity [27], annihilation are critical inputs for - oscillation models by providing more final-state information. More complete models, especially of FSI due to the nucleus, are necessary improve the sensitivity of future searches for - in large liquid argon time projection chambers (LArTPCs) like the Deep Underground Neutrino Experiment (DUNE) [28].
This paper presents the first measurement of antiproton annihilation at rest on argon, performed with the LArIAT liquid argon testbeam experiment [29]. Antiprotons are identified amongst the particles produced in the beamline. A subset of the antiprotons are identified as annihilating at rest and the outgoing particles emerging from the annihilation are compared to a Monte Carlo (MC) simulation. The results presented here represent a first validation of the modeling of nucleon annihilation on argon and may be useful for constraining systematic uncertainties of FSI for pions produced by nucleon annihilation on argon.
@@ II The LArIAT Experiment
The LArIAT experiment collected data at Fermilab’s Test Beam Facility [30] situated in the Meson Center beam line. LArIAT’s primary purpose is measuring charged particle interaction cross sections and developing calibration techniques. The LArIAT TPC operated in conjunction with a series of beamline monitoring subsystems to enable the triggering of detector readouts for individual particles that traverse the entire beamline and interact within the chamber.
@@
Before reaching LArIAT, the secondary beam entered the experimental hall enclosure where it struck a copper target inside a steel collimator to produce a tertiary beam of charged particles. The tertiary beam traveled through a series of beamline instrumentation elements as shown in Fig. 1, which included two time-of-flight (TOF) scintillating paddles and four multi-wire proportional chambers referred to as ‘wire chambers’ (WCs). The trajectory of a charged particle traveling in the beamline was reconstructed by interpolating the straight segments obtained by at least three hits in the upstream WCs in tandem with the downstream ones. This information was used to reconstruct each particle’s three-dimensional trajectory along the tertiary beamline. The polarity of the dipole magnets could be configured to select positively or negatively charged particles. The bending angle and known magnetic field were used to determine the particle’s momentum. In addition, the scintillating paddles measured the particle’s TOF, which was used in tandem with the momentum to determine the particle mass as explained in Sec. III. Detailed discussion on the TOF and WC systems can be found in [29, 31].
Tertiary beamline particles were then directed into the LArTPC, a 170-liter active LAr volume approximately 90 cm in length (, along the beam direction), 47 cm in width (, extending from the anode to the cathode), and 40 cm in height (, pointing upward). The LArTPC and cryostat are shown in Fig. 2. The LArTPC readout comprised two readout-instrumented wire planes composed of 240 wires each (induction and collection), as well as a non-instrumented 225-wire shield plane. For the data used in this measurement, wires were spaced 4 mm apart with a 4 mm separation gap between each of the planes. The induction and collection wire planes were oriented at 60° relative to the beamline direction. Ionization charge drifted along the direction from the cathode to the anode wire planes in a nominal field of 484 V/cm, which resulted in a maximum electron drift period of approximately 320 s. The LArIAT data acquisition system recorded the ionization produced by beamline particles traversing the LArTPC via induced or collected charge on the wire planes, and these signals were digitized with a sampling period of 128 ns.
The analysis presented here uses two weeks of beam-triggered data, taken during LArIAT’s ‘Run II’ data-taking period in 2016, when the tertiary beamline’s dipole magnets were operating in negative-polarity mode in order to steer negatively-charged particles toward the TPC.
@@ III Beamline Reconstruction and Antiproton Selection
All beamline-triggered events are reconstructed to obtain their individual TOF and momentum as they traverse the beamline monitoring systems. It is possible to distinguish different particle types by utilizing the reconstructed TOF and momentum. This is shown in Fig. 3 with overlapping expectations for particles present in the beam, such as deuterons, protons, kaons, pions, muons, and electrons. Horizontal bands present at TOF intervals of 20 ns are caused by multiple high-energy halo muons produced in the secondary beam far upstream of the tertiary beamline. Their time separation is defined by the primary proton beam structure that strike the upstream and downstream TOF scintillator paddles and create a false TOF measurement for particular beam-triggered events.
@@
where is the speed of light and is the length traveled by the particle between the two TOF paddles, 6.65 m. The distribution of the reconstructed mass of beamline particles for Run II is shown in Fig. 4. Particle types are classified according to three distinct regions as shown in the figure:
@@
The beamline mass selection cuts used to identify antiprotons with their reconstructed momentum and TOF are tuned using the Run II positive polarity data where there is a large sample of protons (100,000). Antiprotons are selected with between 800 MeV/c2 and 1100 MeV/c2, inclusively, is chosen to include all of the proton-like mass peak while minimizing contamination from kaons at lower mass and deuterons at higher mass. This cut on isolates events in the (anti)proton-like TOF region. The selection is applied to all the data in Run II taken with -100 A magnet current. The impact of this selection cut is shown in Fig. 3 with the dashed curves outlining the antiproton selection region. Out of the 189,273 events processed in the Run II negative polarity dataset, 520 are selected as having a mass consistent with an antiproton. The range of selected masses is illustrated by the shaded region in Fig. 4.
Kaons could mimic antiproton beamline candidates if they have a short track length due to their similar energy deposition profile. To assess potential kaon contamination in the selected events, a combined fit was applied to the negative polarity data from Fig. 4, with a Gaussian function centered at the fitted kaon mass of 495.2 MeV/c2 and on top of a Landau function to model the non-kaon beam related backgrounds in the calculated mass distribution. The backgrounds fit by the Landau are from the high-energy halo muons striking the TOF paddles. Since the events in these horizontal bands are due to random firings of the trigger from muons produced upstream of the tertiary beamline, they feature a high number of tracks in the TPC with no clear single and unobstructed stopping beam-matched track. These are unlikely to contaminate the sample and are removed during the LArTPC selection stage. Based on an extrapolation of the fit into the proton-mass selection window, the expected contamination from kaons is negligible, with events expected for the full Run II -100 A sample.
@@ IV LArTPC Reconstruction and At-Rest Annihilation Selection
Following the antiproton beamline reconstruction and selection of Sec. III, the second reconstruction and selection phase happens in the LArTPC. In this phase, we use a hybrid manual/semi-automated reconstruction process, culminating in the determination of final candidates on an event-by-event basis using hand-scanning and hand-clustering methods. A ranged-based particle identification (PID) algorithm on both the incoming antiprotons and the outgoing particles from the annihilation vertex.
@@ IV.1 LArTPC Reconstruction & Simulation
LArTPC reconstruction is performed using the LArSoft (version 06.61.00) event processing framework [32, 29, 31] that outputs reconstructed particle objects and their associated kinematics. Particles passing through the LArTPC are reconstructed from ionization signals detected on the wires. These signals are waveforms of ADCs vs. time ticks. Hits within the waveforms are determined by Gaussian fits to over-threshold pulses found in filtered and deconvolved wire waveforms. Hits are typically clustered together by an automated reconstruction algorithm which reconstructs 3D tracks by time-matching clusters of hits across the two 2D wire-planes [33]. LArIAT’s automated reconstruction algorithms were optimized for forward-going particles, and do not reliably reconstruct the more complex, isotropic, multi-track final-states. Unlike other LArTPCs which typically employ three independent wire readout planes, LArIAT has only two planes, which reduces its ability to use the most up-to-date LArTPC reconstruction algorithms. Therefore, to improve its performance, a manual component to the reconstruction is introduced in this analysis.
Contiguous hits that make up the same topological object are manually selected by grouping them together into 2D clusters on each wire plane in a customized event display graphical user interface. These clusters are then used to reconstruct 3D tracks using the standard projection-matching track fitting algorithm in LArSoft [34, 35]. Hits located directly at the annihilation vertex are excluded from clusters associated with the incoming and outgoing particles to avoid complications in the tracking algorithm. This ensures that hits from the stopping antiproton’s Bragg peak, vertex, or from overlapping outgoing ionization trails are not mistakenly included in the wrong clusters. For candidate incoming antiproton tracks, hand-clustering is treated more inclusively by extending the clustering process as close to the annihilation vertex as possible within two reconstructed hits. An example of the results of individual and manual clustering is shown in Fig. 5 where the different clustered objects are individually grouped in different colors and noted from C1 through C5.
@@
In order to gauge the performance of the reconstruction methods, a sample of 1,200 antiprotons entering the LArTPC is simulated. They are generated at the final beamline WC approximately 100 cm upstream of the TPC, with a momentum distribution closely matching what is observed in data by utilizing a Gaussian with similar mean and width. The MC simulation is performed this way to increase the number of annihilations at rest in the active area of the LArTPC, mimicking reality as much as possible. The generated particles decay, interact, and deposit their energy following Geant4 [36] (version 10.3.1) using the QGSP_BERT physics library. To simulate antiproton annihilation on individual nucleons in argon, the Fritiof (FTF) [37, 38] model is used in this paper.
@@ IV.2 LArTPC Particle Identification and Event Selection
A clear, non-overlapping, incoming beamline particle track must come to a stop in the active volume, producing a Bragg peak characterized by a visible increase in charge density along the track (indicated by color in the event display).
@@
Outgoing particles must start at the end of the incoming beamline particle and travel away from the annihilation vertex. These particles are not required to be contained.
@@
The first criterion removes events where the annihilation might have occurred before entering the active region of the LArTPC resulting, in some cases, in an empty event. In data, this criterion additionally removes events containing significant pile-up muon activity which may produce ionization that overlaps and interferes with the actual triggered antiproton track. These pile-ups are caused by additional tracks from muons produced in the upstream beamline elements traveling alongside the true beamline candidates. Since annihilation at rest produces outgoing particles isotropically, which are required to be observed from the second criterion, it is possible that tracks in the LArTPC overlap with the incoming stopping antiproton. The Bragg peak of the selected events ensures that the particle is coming to rest and its track information is passed to the d/d-based PID algorithm to identify the particle type.
@@
The particle associated with a reconstructed track inside the LArTPC is identified by examining the calorimetry information for the track’s reconstructed hits. This is done by looking at the amount of energy deposited per unit length (d/d) along the 3D track object and comparing it to the expected values of d/d for different particle types. In this analysis, particles can be grouped as: proton-like, kaon-like, pion/muon-like, and minimum-ionizing particles without a Bragg peak (non-stopping-like). The PID method consists of calculating a between a track’s modified residual range profile and expected d/d profiles for different particle types. The number of degrees of freedom (NDF) is the number of reconstructed hits used to calculate the total .
To improve the performance of the algorithm, which relies on a well-defined and clear Bragg peak, reconstructed hits at or near the annihilation vertex are omitted, as these can bias the calculation. Specifically, hits within 1 cm of the vertex are excluded as the end of the antiproton Bragg peak can start to merge with the annihilation vertex and overlap with the ionization from outgoing particles. The PID method accounts for this potential under-clustering of the track, which can lead to shorter reconstructed lengths. To mitigate this effect, a correction is applied to the residual range, adjusting it by up to two reconstructed hits—equivalent to 1-2 cm due to the wire-based readout and discretization effects.
In order to determine the selection efficiency, the criteria definition for an “at-rest” annihilation signal at truth-level is defined as an antiproton that annihilates while retaining an amount of kinetic energy less than or comparable to that carried by any one of its constituent quarks in the antiproton’s rest frame, equivalent to 100 MeV due to internal Fermi motion. Since the energy imparted to the final-state annihilation products is the sum of the antiproton’s rest mass energy, its kinetic energy at the moment of interaction, and this random quark motion, this requirement on the true signal ensures events whose final-state metrics (energy and track multiplicity) are comparable to true at-rest annihilation within 5-10%.
@@
After visually scanning 1,200 simulated antiproton events, 80 stopping candidates were selected. Of these candidates selected through hand-scanning, 69 truly annihilate at rest, 9 retain some residual kinetic energy less than 60 MeV, while the remaining 2 annihilate with nearly 200 MeV of remaining energy. Based on the signal definition above, the mis-ID rate for visual inspection is therefore estimated to be %. Of the initial 1,200 events, 140 are considered signal as they have true final kinetic energy of less than 100 MeV. This selection of antiproton annihilation at rest events results in a selection efficiency of 56% and a purity of 98%.
@@
All of the 80 MC signal candidates are also processed through the PID algorithm. The results indicate that 70 of the 80 events are identified as proton-like, 6 as kaon-like, and 4 as stopping muon/pion-like. To evaluate the performance of the PID algorithm, the d/d distributions for the three categories of antiproton candidates are compared with the expected values, as shown in top of Fig. 7. The proton-like candidates closely align with the proton expectation curve. In comparison, for proton tracks that interact before coming to a full stop, or have a mis-reconstructed endpoint that is placed upstream of the true endpoint, the d/d profile tends to more closely resemble a kaon-like hypothesis. The stopping muon/pion-like events, which correspond well with the pion/muon expectation curve, are rejected from our sample. The two antiproton candidates that have over 200 MeV of kinetic energy were identified as 1 kaon-like and 1 muon/pion-like and would be removed at this stage. Based on these observations, all 10 non-proton-like candidates are excluded from the sample, resulting in a PID efficiency of 88%.
@@
Following the established selection criteria, the number of data events from the data period described in Sec. II is reduced from 520 initial candidates in the selected mass range, down to 11 candidate stopping antiprotons annihilating at rest. The reconstructed beamline mass distribution shown in Fig. 6 with a fitted Gaussian centered at 933.2 27.0 MeV/c2 suggests that these are all true antiprotons. All 11 candidates are also identified as proton-like using the PID calculation. The d/d as a function of the residual range distribution of the antiproton candidates is shown in bottom of Fig. 7. All 11 candidates align well with the proton-like expected curve confirming that they are expected to be true antiproton annihilations at rest given the small misidentification rate of % evaluated from the simulation.
A summary of the effect of each selection cut is shown in Table 1. High energy muon pile-ups are not simulated, so a difference in the relative number of events removed from data and simulation is expected for the second cut. A collection of selected antiproton annihilation event displays from data and simulation is provided in Appendix A.